jamal is 12 years older than twice stephanie's age. the sum of their ages is 45.

Answer by stanbon(75887)

jamal is 12 years older than twice stephanie's age. the sum of their ages is 45.
------
Equation:
j + s = 45
j = 2s+12
-----
Substitute for "j" and solve for "s":
2s+12 + s = 45
3s = 33
s = 11 (Stephanie's age now)
-----
Solve for "j":
j + s = 45
j + 11 = 45
j = 34 (Jamal's age now)
==============
Cheers,
Stan H.
==============